# Robust Cooperative Multi-Agent Reinforcement Learning: A Mean-Field Type Game Perspective

### v1

@@ 3 Robust Mean-Field Control
$J^{\gamma}(\mathcal{u}^{1},\mathcal{u}^{2})=$ $\mathbb{E}\sum_{t=0}^{T}\Big{[}\lVert x_{t}-\bar{x}_{t}\rVert^{2}_{Q_{t}}+\lVert\bar{x}_{t}\rVert^{2}_{\bar{Q}_{t}}+\lVert u^{1}_{t}-\bar{u}^{1}_{t}\rVert^{2}+\lVert\bar{u}^{1}_{t}\rVert^{2}-\gamma^{2}\big{(}\lVert u^{2}_{t}-\bar{u}^{2}_{t}\rVert^{2}+\lVert\bar{u}^{2}_{t}\rVert^{2}\big{)}$ (8)
@@
$\min_{K^{1}_{t},L^{1}_{t}}\max_{K^{2}_{t},L^{2}_{t}}J^{\gamma}((K^{1}_{t},K^{2}_{t}),(L^{1}_{t},L^{2}_{t}))=\min_{K^{1}_{t}}\max_{K^{2}_{t}}J^{\gamma}_{y}(K)+\min_{L^{1}_{t}}\max_{L^{2}_{t}}J^{\gamma}_{z}(L)$
@@
The proof of this result can be found in the Supplementary Materials. The above theorem states that, if for a given $\gamma$, conditions (12) and (16) are satisfied (given that $M^{\gamma}_{t}$ and $\bar{M}^{\gamma}_{t}$ are defined by (13)), then not only is $\gamma$ a viable attenuation level for the original Robust multi-agent control problem (1)-(4), but the Nash equilibrium for the ZS-MFTG also yields the robust controller $u^{1,i*}_{t}=-K^{1*}_{t}(x^{i}_{t}-\bar{x}_{t})-L^{1*}_{t}\bar{x}_{t}$ for the original finite-agent game. Condition (16) is strictly stronger than condition (15) but approaches (16) as $N\rightarrow\infty$.
@@ 4 Reinforcement Learning for Robust Mean-Field Control
Approximation of mean-field terms using $M$ agents: Since simulating infinitely many agents is impractical, in this section we outline how to use a set of $2\leq M<\infty$ agents to approximately simulate the mean-field in a MFTG. Each of the $M$ agents has state $x^{i}_{t}$ at time $t$ where $i\in[M]$. The agents follow controllers linear in their private state and empirical mean-field, $x^{i}_{t}$ and $\tilde{z}_{t}$, respectively:
$u^{1}_{t}=-K^{1}_{t}(x^{i}_{t}-\tilde{z}_{t})-L^{1}_{t}\tilde{z}_{t},\hskip 14.22636ptu^{2}_{t}=K^{2}_{t}(x^{i}_{t}-\tilde{z}_{t})+L^{2}_{t}\tilde{z}_{t},$
where the empirical mean-field is $\tilde{z}_{t}:=\frac{1}{M}\sum_{i=1}^{M}x^{i}_{t}$. Under these control laws, the dynamics of agent $i\in[M]$ are
@@
and the dynamics of the empirical mean-field $\tilde{z}_{t}$ is
[…]
The cost of each agent is
@@
Now, similarly to the previous section, we define $y^{i}_{t}=x^{i}_{t}-\tilde{z}_{t}$. The dynamics of $y^{i}_{t}$ are
$y^{i}_{t+1}$ $=(A_{t}-B_{t}K^{1}_{t}+K^{2}_{t})y^{i}_{t}+\tilde{\omega}^{i}_{t+1},\hskip 5.69046pt\text{where}\hskip 5.69046pt\tilde{\omega}^{i}_{t+1}=\frac{M-1}{M}\omega^{i}_{t+1}-\frac{1}{M}\sum_{j\neq i}\omega^{j}_{t+1}.$
The cost can then be decomposed in a manner similar to (11):
@@
Closed form expressions of the bounds can be found in the proof given in the Supplementary materials. The linear rate of convergence is made possible by building upon the convergence analysis of descent ascent in (Fallah et al., 2020) due to the convex-concave nature of the cost function (18). The proof generalizes the techniques used in (Fallah et al., 2020) to stochastic unbiased gradients by utilizing the fact that the bias in stochastic gradients ~$\tilde{\nabla}_{j}\tilde{J}^{i,\gamma}_{t}$ for $j\in\{1,2\}$ can be reduced by reducing the smoothing radius $r$. This in turn causes an increase in the variance of the stochastic gradient which is controlled by increasing the mini-batch size $N_{b}$.
Now we present the non-asymptotic convergence guarantee of the paper stating that even though each iteration of the outer loop (as timestep $t$ moves backwards-in-time) accumulates error, if the error in each outer loop iteration is small enough, the total accumulated error will also be small enough. The proof can be found in the Supplementary Materials.
@@ 6 Conclusion
In this paper, we solve an MARL problem with the objective of designing robust controllers in the presence of modeled and un-modeled uncertainties. We introduce the concept of Robust Mean Field Control (RMFC) problem as the limiting problem when the number of agents grows to infinity. We then establish a connection with Zero-Sum Mean-Field Type Games (ZS-MFTG). We resort to the Linear-Quadratic (LQ) structure which, combined with the mean-field approximation, helps to have a more tractable model and to help resolve the analytical difficulty induced by the distributed information structure. This helps us obtain sufficient conditions for robustness of the problem as well as characterization of the robust control policy. We design and provide non-asymptotic analysis of a receding-horizon based RL algorithm which renders the non-coercive cost as convex-concave. Through numerical analysis the receding-horizon approach is shown to ameliorate the overshooting problem observed in the performance of the vanilla algorithm.
In future work we would like to explore this type of robust mean-field problems beyond the LQ setting and to develop RL algorithms which go beyond the gradient descent-ascent updates used in this paper. Furthermore, our work is a first step in the direction of using mean-field approximations to study robust MARL problems which occur in many real-world scenarios, but the study of concrete examples is left for future work. [^2] Disclaimer: This paper was prepared for informational purposes in part by the Artificial Intelligence Research group of JP Morgan Chase & Co and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.

### v2

@@ 3 Robust Mean-Field Control
$J^{\gamma}(\mathcal{u}^{1},\mathcal{u}^{2})=$ $\mathbb{E}\sum_{t=0}^{T-1}\Big{[}\lVert x_{t}-\bar{x}_{t}\rVert^{2}_{Q_{t}}+\lVert\bar{x}_{t}\rVert^{2}_{\bar{Q}_{t}}+\lVert u^{1}_{t}-\bar{u}^{1}_{t}\rVert^{2}+\lVert\bar{u}^{1}_{t}\rVert^{2}-\gamma^{2}\big{(}\lVert u^{2}_{t}-\bar{u}^{2}_{t}\rVert^{2}+\lVert\bar{u}^{2}_{t}\rVert^{2}\big{)}$ (8)
@@
$\min_{K^{1},L^{1}}\max_{K^{2},L^{2}}J^{\gamma}((K^{1},K^{2}),(L^{1},L^{2}))=\min_{K^{1}}\max_{K^{2}}J^{\gamma}_{y}(K)+\min_{L^{1}}\max_{L^{2}}J^{\gamma}_{z}(L)$
@@
The proof of this result can be found in the full version of this paper (Zaman et al., 2024). The above theorem states that, if for a given $\gamma$, conditions (12) and (16) are satisfied (given that $M^{\gamma}_{t}$ and $\bar{M}^{\gamma}_{t}$ are defined by (13)), then not only is $\gamma$ a viable attenuation level for the original Robust multi-agent control problem (1)-(4), but the Nash equilibrium for the ZS-MFTG also yields the robust controller $u^{1,i*}_{t}=-K^{1*}_{t}(x^{i}_{t}-\bar{x}_{t})-L^{1*}_{t}\bar{x}_{t}$ for the original finite-agent game. Condition (16) is strictly stronger than condition (15) but approaches (16) as $N\rightarrow\infty$.
@@ 4 Reinforcement Learning for Robust Mean-Field Control
Approximation of mean-field terms using $M$ agents: Since simulating infinitely many agents is impractical, in this section we outline how to use a set of $2\leq M<\infty$ agents to approximately simulate the mean-field in a MFTG. Each of the $M$ agents has state $x^{i}_{t}$ at time $t$ where $i\in[M]$. The agents follow controllers linear in their private state and empirical mean-field, $x^{i}_{t}$ and $\tilde{z}_{t}$, respectively, $u^{1}_{t}=-K^{1}_{t}(x^{i}_{t}-\tilde{z}_{t})-L^{1}_{t}\tilde{z}_{t},u^{2}_{t}=K^{2}_{t}(x^{i}_{t}-\tilde{z}_{t})+L^{2}_{t}\tilde{z}_{t},$ where the empirical mean-field is $\tilde{z}_{t}:=\frac{1}{M}\sum_{i=1}^{M}x^{i}_{t}$. Under these control laws, the dynamics of agent $i\in[M]$ are
@@
and the dynamics of the empirical mean-field $\tilde{z}_{t}$ is $\tilde{z}_{t+1}=(\tilde{A}_{t}-\tilde{B}_{t}L^{1}_{t}+L^{2}_{t})\tilde{z}_{t}+\tilde{\omega}^{0}_{t+1},\hskip 5.69046pt$ where $\tilde{\omega}^{0}_{t+1}=\bar{\omega}_{t}+\frac{1}{M}\sum_{i=1}^{M}\omega^{i}_{t+1}.$ The cost of each agent is
@@
Now, similarly to the previous section, we define $y^{i}_{t}=x^{i}_{t}-\tilde{z}_{t}$. The dynamics of $y^{i}_{t}$ are $y^{i}_{t+1}=(A_{t}-B_{t}K^{1}_{t}+K^{2}_{t})y^{i}_{t}+\tilde{\omega}^{i}_{t+1},$ where $\tilde{\omega}^{i}_{t+1}=\frac{M-1}{M}\omega^{i}_{t+1}-\frac{1}{M}\sum_{j\neq i}\omega^{j}_{t+1}$. The cost can then be decomposed in a manner similar to (11):
@@
Closed form expressions of the bounds can be found in the proof given in the full version of the paper (Zaman et al., 2024). The linear rate of convergence is made possible by building upon the convergence analysis of descent ascent in (Fallah et al., 2020) due to the convex-concave nature of the cost function (18). The proof generalizes the techniques used in (Fallah et al., 2020) to stochastic unbiased gradients by utilizing the fact that the bias in stochastic gradients ~$\tilde{\nabla}_{j}\tilde{J}^{i,\gamma}_{t}$ for $j\in\{1,2\}$ can be reduced by reducing the smoothing radius $r$. This in turn causes an increase in the variance of the stochastic gradient which is controlled by increasing the mini-batch size $N_{b}$.
Now we present the non-asymptotic convergence guarantee of the paper stating that even though each iteration of the outer loop (as timestep $t$ moves backwards-in-time) accumulates error, if the error in each outer loop iteration is small enough, the total accumulated error will also be small enough. The proof can be found in the complete version of the paper (Zaman et al., 2024).
@@ 6 Conclusion
In this paper, we solve an MARL problem with the objective of designing robust controllers in the presence of modeled and un-modeled uncertainties. We introduce the concept of Robust Mean Field Control (RMFC) problem as the limiting problem when the number of agents grows to infinity. We then establish a connection with Zero-Sum Mean-Field Type Games (ZS-MFTG). We resort to the Linear-Quadratic (LQ) structure which, combined with the mean-field approximation, helps to have a more tractable model and to help resolve the analytical difficulty induced by the distributed information structure. This helps us obtain sufficient conditions for robustness of the problem as well as characterization of the robust control policy. We design and provide non-asymptotic analysis of a receding-horizon based RL algorithm which renders the non-coercive cost as convex-concave. Through numerical analysis the receding-horizon approach is shown to ameliorate the overshooting problem observed in the performance of the vanilla algorithm. In future work we would like to explore this type of robust mean-field problems beyond the LQ setting and to develop RL algorithms which go beyond the gradient descent-ascent updates used in this paper. Furthermore, our work is a first step in the direction of using mean-field approximations to study robust MARL problems which occur in many real-world scenarios, but the study of concrete examples is left for future work.
@@ Acknowledgments
Disclaimer:
This paper was prepared for informational purposes in part by the Artificial Intelligence Research group of JP Morgan Chase & Co and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.